Rylee88:

To kick off the summer, Leah is planning a "School's Out!" party for all her friends. She and her mom pick up a giant sub and a vat of root beer from Subs & Suds. When they get to the park, they cut the sub into 18 equal sections. Each section is 1/3 of a foot long. Which equation can you use to find the length g of the giant sub? g+18= 1/3 18g=1/33 g18=1/3 g–18=1/3 Solve this equation for g to find the length of the giant sub. To write a fraction, use a slash ( / ) to separate the numerator and denominator. feet

SmokeyBrown:

If there are 18 sections of the sub, and each section is 1/3 of a foot long, then the total length of the sub should be g = 18 * 1/3 Can you see that any of the answer choices would be equivalent to this equation?

nightshade0773:

IM A SHE NOT HE

Rylee88:

OOOOOo

Rylee88:

My bad

Alexis1415:

nightshade answer the question....lol

nightshade0773:

lol

Alexis1415:

@smokeybrown wrote:
If there are 18 sections of the sub, and each section is 1/3 of a foot long, then the total length of the sub should be g = 18 * 1/3 Can you see that any of the answer choices would be equivalent to this equation?
answer this nightshade

nightshade0773:

g18=1/3

SmokeyBrown:

I see that one of the answers "g18 = 1/3" is very similar to "g/18 = 1/3", which would be correct. Is it possible that one of the "/" was missing from that somehow?

Alexis1415:

??

Alexis1415:

shade.